With the Haber process how many grams of hydrogen would be needed to react with 13g of N2


1
Expert's answer
2021-09-07T02:08:18-0400

N2 + 3 H2 = 2 NH3

n (H2) = 3 x n (N2)

n = m / M

M (N2) = 28 g/mol

M (H2) = 2 g/mol

n (N2) = 13 / 28 = 0.46 mol

n (H2) = 3 x n (N2) = 3 x 0.46 = 1.4 mol

m (H2) = n x M = 1.4 x 2 = 2.8 g